a solution of H2SO4 with a molal concentration of 3.58m has a density of 1.200g/mL. what is the molar concentration of this olution?
1
Expert's answer
2019-09-25T05:44:20-0400

Molar concentration of 3.58m means there are 3.58m of H2SO4 in 1 kg of solvent (assuming its water).

Mass=molar concentration*molecular weight

Mass (of H2SO4) = 3.58 * 98.079= 351.12 g

Mass of solution= 351.12 (H2SO4) + 1000 (H2O) = 1351 g

Volume = mass / density

Volume = 1351 / 1.200 * 1 = 1125 mL

Molar concentration = mol in L

Molar concentration = 3.58 / 1.125 = 3.18 M

3.18M